# The three phases of self-gravitating scalar field ground states

### v1

@@ Introduction
Most work on ULDM assumes that each dark matter halo has a solitonic core at its center.In ULDM, solitons condense quickly enough that most halos would contain a soliton Levkov et al. (2018); Eggemeier and Niemeyer (2019); Kirkpatrick et al. (2020); Chen et al. (2021); Hertzberg et al. (2020); Kirkpatrick et al. (2022); García et al. (2023). The role solitons could play as a solution to the cusp-core problem contributed to ULDM rising in popularity as a dark matter candidate Su and Chen (2011); Hui et al. (2017). While recent work has investigated the relative size and shape of solitonic cores in simulated ULDM halos Chan et al. (2022); Yavetz et al. (2022); Chan et al. (2022); Kendall et al. (2023); Zagorac et al. (2023), their existence in centers of halos has not been called into question. Nevertheless, the string axiverse generically predicts the existence of multiple ultralight fields with nonzero self- and interspecies interactions Arvanitaki et al. (2010); Arvanitaki and Dubovsky (2011), which in turn impacts the core profiles of such structures Glennon et al. (2023). Therefore, the mapping from interaction strengths to core profiles to the diversity of halos found in nature could enable constraints on axiverse models.
These results also have implications for inflation. In some inflationary models, the reheating epoch is analogous to structure formation in ULDM  Easther et al. (2011); Musoke et al. (2020), going as far as the formation of inflaton halos and solitonic inflaton stars Eggemeier and Niemeyer (2019). Interactions between multiple scalar inflation fields would significantly change these collapsed states, impacting their dynamics during reheating and the consequent observational predictions Hotinli et al. (2018).
@@
Numerical results for phase transitions
@@ Appendix C Imaginary Time Solver
We find spherically symmetric ground states by solving the spherically symmetric versions of the GPP equations of motion in imaginary time. The steps in our imaginary time evolution obey a “kick-drift-kick” algorithm (as in Edwards et al. (2018); Glennon and Prescod-Weinstein (2021); Padmanabhan et al. (2020); Musoke (2024)). Schematically, this can be represented for $n$ iterations of timestep $h$ for wavefunction $i$ as:
@@
$V_{\rm{eff}}$ $=m_{i}\Phi+\frac{1}{2}\sum_{j}m_{j}\lambda_{ij}\left|\psi_{j}\right|^{2}\psi_{i}\,.$ (29)
@@
The time evolution in nSPIRal conserves the total mass in each field and introduces a hard boundary at a given maximum radial distance $r=r_{\text{max}}$. The boundary value of each field is fixed such that $\left.r\psi_{j}\right|_{r=0}=\left.r\psi_{j}\right|_{r_{\text{max}}}=0$; the latter is equivalent to specifying $\psi_{j}(r_{\text{max}})=0$. We therefore take care to use only large enough $r_{\rm{max}}$ such that $\psi_{j}(r_{\rm{max}})\rightarrow 0$ for each field. In addition to $r_{\rm{max}}$, we must specify our grid size $n_{g}$, number of timesteps $n$, and initial conditions for the $\psi_{j}$ profiles.

### v2

@@ Introduction
Most work on ULDM assumes that each dark matter halo has a solitonic core at its center.In ULDM, solitons condense quickly enough that most halos would contain a soliton Levkov et al. (2018); Eggemeier and Niemeyer (2019); Kirkpatrick et al. (2020); Chen et al. (2021); Hertzberg et al. (2020); Kirkpatrick et al. (2022); García et al. (2023). The role solitons could play as a solution to the cusp-core problem contributed to ULDM rising in popularity as a dark matter candidate Su and Chen (2011); Hui et al. (2017). While recent work has investigated the relative size and shape of solitonic cores in simulated ULDM halos Chan et al. (2022); Yavetz et al. (2022); Chan et al. (2022); Kendall et al. (2023); Zagorac et al. (2023); Dave and Goswami (2023, 2024); Chakrabarti et al. (2022), their existence in centers of halos has not been called into question. Nevertheless, the string axiverse generically predicts the existence of multiple ultralight fields with nonzero self- and interspecies interactions Arvanitaki et al. (2010); Arvanitaki and Dubovsky (2011), which in turn impacts the core profiles of such structures Glennon et al. (2023). Therefore, the mapping from interaction strengths to core profiles to the diversity of halos found in nature could enable constraints on axiverse models.
These results also have implications for inflation. In some inflationary models, the reheating epoch is analogous to structure formation in ULDM  Easther et al. (2011); Musoke et al. (2020), going as far as the formation of inflaton halos and solitonic inflaton stars Eggemeier and Niemeyer (2019). Interactions between multiple scalar inflaton fields would significantly change these collapsed states, impacting their dynamics during reheating and the consequent observational predictions Hotinli et al. (2018).
@@
Numerical results and phase diagram
@@ Appendix C Imaginary Time Solver
We find spherically symmetric ground states by solving the spherically symmetric versions of the GPP equations of motion in imaginary time and using the code units described above. The steps in our imaginary time evolution obey a “kick-drift-kick” algorithm (as in Edwards et al. (2018); Glennon and Prescod-Weinstein (2021); Padmanabhan et al. (2020); Musoke (2024)). Schematically, this can be represented for $n$ iterations of timestep $h$ for wavefunction $i$ as:
@@
$V_{\rm{eff}}$ $=m_{i}\Phi+\frac{1}{4}\sum_{j}m_{j}\lambda_{ij}\left|\psi_{j}\right|^{2}\psi_{i}\,.$ (29)
@@
  4. 4.
    re-normalizing $\psi_{i}$ so that the total mass in each field is conserved.
Additionally, nSPIRal introduces a hard boundary at a given maximum radial distance $r=r_{\text{max}}$. The boundary value of each field is fixed such that $\left.r\psi_{j}\right|_{r=0}=\left.r\psi_{j}\right|_{r_{\text{max}}}=0$; the latter is equivalent to specifying $\psi_{j}(r_{\text{max}})=0$. We therefore take care to use only large enough $r_{\rm{max}}$ such that $\psi_{j}(r_{\rm{max}})\rightarrow 0$ for each field. In addition to $r_{\rm{max}}$, we must specify our grid size $n_{g}$, number of timesteps $n$, and initial conditions for the $\psi_{j}$ profiles.